# Beyond the Black Box: Theory and Mechanism of Large Language Models

### v1

@@ 6.2.2 In-Context Learning
Empirical Camp. This camp directly explores and investigates the characteristics of the ICL process in large models from experiments rather than theory, providing empirical insights for theoretical analysis of ICL. icl_Garg_casestudy examine the ability of Transformers to be trained on well-defined tasks, such as linear tasks, and ultimately learn context. It has been found that the Transformer can achieve predictive performance comparable to least squares algorithm. As the problem becomes sparse, the prediction error of in-context learning (ICL) will be comparable to the solution of the Lasso problem. Additionally, they investigate more complicated tasks, such as two-layer neural networks and four-layer decision trees, and find that Transformers could effectively learn and generalize on these function classes as well. In contrast to simulation setting in icl_Garg_casestudy, icl_diff_setting conduct extensive exploration using a series of LLMs, including GPT-3, InstructGPT, Codex, PaLM, and Flan-PaLM, across different configurations. Firstly, they examined the ICL setting with flipped labels to assess the models’ ability to override prior knowledge. They note that smaller models primarily rely on semantic priors from pretraining during ICL inference, thus often disregarding label flips in the context. Conversely, larger models, despite having stronger semantic priors, demonstrate the capability to override these priors when faced with label flips. Further, they investigate the ICL setting with semantically unrelated labels, highlighting that sufficiently large models can perform linear classification tasks under this setting. In addition, they evaluate models fine-tuned with instructions and find that instruction tuning notably enhanced the utilization of semantic priors compared to learning input-label mappings from contextual demonstrations. Another influential work is the study by emnlp_bestpaper on the mechanism of ICL from the perspective of information flow, which find that in input-label pairs, label tokens act as anchors. Initially, the semantic information from the context aggregates into the token representation of the label tokens at the shallower layers of LLMs, and then the final predictions of LLMs reference the aggregated information in the label tokens. Building on this finding, anchor-based re-weighting methods, demonstration compression techniques, and diagnostic analysis frameworks for ICL errors are further proposed, yielding the expected results and validate the analysis.
@@ 6.2.3 Inference-Time Scaling
Inference-time scaling represents a fundamental shift in the deployment of LLMs, where the reasoning capacity is no longer viewed as a static property of the model’s parameters but as a dynamic function of the computational resources allocated during interaction [chen2025towards, snell2024scaling]. This paradigm is primarily established through the Chain-of-Thought (CoT) mechanism and various external search-based algorithms that extend the model’s “thinking” process [wei2022chain, yao2024tree, kang2024mindstar, zhang2024rest, feng2023alphazero]. Based on current research, we categorize the theoretical investigations of this phenomenon into the following three core dimensions.
@@
Theoretical Expressivity and Boundaries of CoT. A foundational line of inquiry examines how the introduction of intermediate reasoning steps alters the inherent computational limits of the Transformer architecture. Theoretical analysis suggests that CoT serves as an effective depth-extender for auto-regressive models. feng2023towards utilize circuit complexity theory to prove that finite-depth Transformers can perfectly execute these tasks by extending their effective depth linearly with the number of generated reasoning steps. This is further formalized by li2024chain, which demonstrates that while constant-depth Transformers without CoT are restricted to parallelizable complexity classes such as $AC^{0}$ or $TC^{0}$, the addition of reasoning steps enables the model to solve any problem within the $P/poly$ complexity class. To bridge these theoretical findings with practical performance, chen2024unlocking introduce the Reasoning Boundary Framework (RBF) to define the quantitative limits of model performance across different task complexities. Furtherly, sprague2025to reveal that CoT benefits are predominantly concentrated in mathematical and symbolic tasks, providing minimal gains in general knowledge retrieval or tasks lacking explicit logical operators.
@@
Mechanistic Origins and Internal Dynamics. Understanding how these reasoning capabilities emerge and are organized internally is critical for piercing the “black box” of LLM intelligence. dutta2024how identify a functional bifurcation within the Transformer layers: lower layers primarily transform representations from pre-training priors to context-aware embeddings, while middle-to-higher layers act as answer writers that causally integrate information from previously generated CoT steps. Alternatively, 10.5555/3737916.3740933 discovers that reasoning circuits only form through “grokking”, when training significantly beyond the point of overfitting, allowing for robust out-of-distribution generalization that shallow statistical matching cannot achieve. Furthermore, li2025training provide a convergence analysis for how gradient descent optimization enables non-linear Transformers to learn CoT, quantifying the sample complexity required to remain robust against noisy context examples. Beyond explicit prompting, wang2024chain demonstrates that reasoning trajectories are an intrinsic capability of pre-trained models that can be elicited through specialized decoding strategies, such as exploring top-$k$ alternative tokens to find valid reasoning paths without human instructions.
@@
Reliability & Generalization Limits. Despite the empirical success of inference-time scaling, researchers have identified significant bottlenecks related to error accumulation and distributional sensitivity. zhao2025chainofthoughtreasoningllmsmirage argue that the efficacy of CoT is inherently fragile, relying heavily on the consistency between the training reasoning paths and the test-time queries, which suggests that models may be performing advanced pattern matching rather than deep logical deduction. This fragility is particularly evident in complex environments. stechly2024chain show that CoT performance degrades rapidly when task scale or complexity exceeds the scope of the provided examples. To mitigate the “snowball error” effect, where a single early mistake leads to catastrophic reasoning failure, pmlr-v267-gan25a demonstrate that external scaling through search algorithms like Best-of-N and Monte Carlo Tree Search (MCTS) effectively expands the solution space and allows verifiers to select correct paths. To evaluate the quality of these dynamic steps, pmlr-v267-ton25a propose an “Information-Gain” metric, identifying “thinking tokens” that significantly reduce the predictive cross-entropy loss of the final answer, thereby providing a principled tool for diagnosing and optimizing the reasoning process.
@@ 6.3.1 The Overthinking Phenomenon
The empirical success of inference-time scaling, exemplified by leading reasoning models [openai2024reasoning, guo2025deepseek], has popularized the paradigm that more computation leads to better reasoning. However, recent research has identified a critical counter-phenomenon known as “overthinking”, where models generate excessive, redundant, or even erroneous reasoning steps for tasks that are either inherently simple or unsolvable.
Traditional intuition suggests that longer CoT sequences facilitate better task decomposition. However, wu2025lessunderstandingchainofthoughtlength challenge this assumption by demonstrating an inverted U-shaped relationship between reasoning length and accuracy. This work posits that while length helps reduce sub-task difficulty, it simultaneously increases the risk of error accumulation. This balance is further discussed by gan2025cotspacetheoreticalframeworkinternal, which treats CoT as an optimization process in continuous semantic space, identifying a fundamental trade-off between “under-reasoning” (underfitting) and “overthinking” (overfitting).
Overthinking is particularly prevalent when models mimic long-reasoning behaviors for trivial queries. pmlr-v267-chen25bx observe that models often generate verbose reasoning for extremely simple arithmetic, significantly increasing latency and cost without any performance gain. This dilemma is also explored in agentic contexts, cuadron2025dangeroverthinkingexaminingreasoningaction highlight how excessive searching and value modeling can hinder the efficiency of logical agents. A more severe form of overthinking occurs when models encounter pathological queries. fan2025missing find that models optimized for reasoning tend to fall into redundant loops of self-doubt and hallucination when faced with unsolvable problems due to missing premises. This behavior is attributed to current RL mechanisms that over-reward detailed CoT. To address these inefficiencies, li2025compressingchainofthoughtllmsstep propose information-theoretic metrics to quantify the information contribution of each reasoning step. Their findings suggest that a significant portion of steps in modern reasoning models are low-entropy redundancies that can be compressed without compromising accuracy.
@@ 6.3.2 Latent Reasoning
Recent research explores bypassing these limitations by conducting reasoning directly within the model’s latent activation space. hao2025traininglargelanguagemodels propose COCONUT, which allows models to reason in a continuous latent space by feeding hidden states back as subsequent inputs, enabling the encoding of multiple reasoning branches simultaneously. The architectural backbone of latent reasoning often involves weight-tied recurrence. saunshi2025reasoninglatentthoughtspower posit that reasoning performance is primarily driven by computational depth rather than total parameters. The study demonstrates that looped architectures can simulate CoT internally through “latent thoughts”. These models show strong inductive biases for reasoning tasks, suggesting that latent iterations can efficiently substitute for explicit token generation.
A pivotal development in this area is the study of the superposition mechanism within continuous latent spaces. zhu2025reasoningsuperpositiontheoreticalperspective demonstrate that the model can maintain multiple reasoning trajectories in a state of superposition within the continuous latent space, facilitating implicit parallel thinking that far exceeds the capabilities of traditional serial reasoning. The emergence of this mechanism is deeply tied to the training dynamics. zhu2025emergencesuperpositionunveilingtraining further characterize it as a two-stage process and elucidate how the model can simultaneously maintain multiple inference traces in a continuous latent space, thereby achieving implicit parallel thinking. Despite its efficiency, latent reasoning introduces unique challenges. xu2025formalcomparisonchainofthoughtlatent highlight that while latent thoughts support more efficient parallel computation, discrete CoT remains superior for tasks requiring stochastic decoding to approximate complex solutions.
@@ 7.1 Fundamental Problems
The evaluation deeply intertwines with theoretical questions of metrology and security. Unlike in traditional machine learning, where concepts like “robustness”, “fairness”, and “privacy” were often well-defined and could be formalized using precise mathematical objectives and metrics, the current landscape of LLMs presents a new challenge [chang2024survey, anwar2024foundational, dominguez-olmedo2025training, hardt2025performative]. The core, fundamental problems in the Evaluation Stage are therefore:
@@
These two fundamental problems define the ultimate theoretical challenges in the Evaluation Stage. In fig. 7 we provide a landscape of the corresponding topics. To begin answering these profound questions, the academic community has initiated several concrete lines of research, each tackling a specific evaluation tool or observable phenomenon. In what follows, we will review these research efforts, detailing how the study of specific strategies provides valuable insights.
@@ 7.2.1 Benchmark Theory
Benchmark Validity. A growing body of theoretical and empirical work suggests that high performance on static benchmarks may not correlate with true, generalized capabilities [zhang2025benchmark]. Several key limitations in current benchmark-based evaluations are identified. One primary issue is “Shortcut Learning” [du2023shortcut], where models are found to rely on spurious, non-robust features or biases within a dataset rather than mastering the high-level semantic or reasoning skills the benchmark purports to measure [chandak2025eliminating]. This leads to a fundamental lack of robustness and exposes that models may be overfitting to the specific artifacts of the test set rather than the underlying task [lunardi2025robustness]. Furthermore, many traditional benchmarks are facing issues of saturation, where top-tier models approach perfect scores, limiting the benchmark’s ability to distinguish between SOTA models. To address this, zhou2025lost have proposed applying frameworks from psychometrics, such as Item Response Theory, to analyze benchmark quality. This has revealed that many benchmarks suffer from a low “difficulty ceiling” and “item saturation”. zhang2024inherent further derive analysis from the perspective of social choice theory, and demonstrate the trade-offs between benchmark diversity and stability. Finally, the single-scalar scores produced by most benchmarks obscure the complex combination of skills required for a task. kim2025benchmark aim to mechanistically diagnose benchmark composition, decomposing performance into contributions from discrete cognitive abilities.
@@
LLM-as-a-Judge. To overcome the limitations of static benchmarks, especially for evaluating open-ended generation, the “LLM-as-a-Judge” (LLM-Judges) paradigm has become widespread [gu2025surveyllmasajudge]. This approach leverages a powerful LLM to score or rank the outputs of other models. However, this has shifted the theoretical burden from evaluating task performance to evaluating the evaluator itself. This paradigm rests on several core assumptions that LLMs can serve as valid human proxies, are capable evaluators, are scalable, and are cost-effective, unfortunately, all of which are being theoretically challenged [dorner2025limits]. Researchers have argued that LLM-Judges may possess only “face validity” rather than true, robust evaluative capacity [chehbouni2025neither]. Two critical flaws being investigated are low reliability and poor psychometric validity. The common practice of using fixed randomness to ensure reproducibility does not guarantee internal consistency. By applying psychometric measures like McDonald’s omega and repeating evaluations with different random seeds, studies have found the internal consistency reliability of LLM-Judges to be questionable [schroeder2024can]. Also, LLM-Judges suffer from severe design flaws that can render their judgments noisy. Key issues include low “schematic adherence” and “factor collapse” causing the misalignment between the evaluation results and criteria [feuer2025judgment]. Finally, LLM-Judges have been shown to exhibit numerous systematic biases, including position bias, verbosity bias, and authority bias [ye2024justice, chen-etal-2024-humans]. While some work suggests these biases can be partially mitigated through robust prompting with detailed scoring rubrics [gao2025evaluating], the theoretical understanding of these biases remains a critical open area.

### v2

@@ 6.2.2 In-Context Learning
Empirical Camp. This camp directly explores and investigates the characteristics of the ICL process in large models from experiments rather than theory, providing empirical insights for theoretical analysis of ICL. Garg et al. (2022) examine the ability of Transformers to be trained on well-defined tasks, such as linear tasks, and ultimately learn context. It has been found that the Transformer can achieve predictive performance comparable to least squares algorithm. As the problem becomes sparse, the prediction error of in-context learning (ICL) will be comparable to the solution of the Lasso problem. Additionally, they investigate more complicated tasks, such as two-layer neural networks and four-layer decision trees, and find that Transformers could effectively learn and generalize on these function classes as well. In contrast to simulation setting in Garg et al. (2022), Wei et al. (2023) conduct extensive exploration using a series of LLMs, including GPT-3, InstructGPT, Codex, PaLM, and Flan-PaLM, across different configurations. Firstly, they examined the ICL setting with flipped labels to assess the models’ ability to override prior knowledge. They note that smaller models primarily rely on semantic priors from pretraining during ICL inference, thus often disregarding label flips in the context. Conversely, larger models, despite having stronger semantic priors, demonstrate the capability to override these priors when faced with label flips. Further, they investigate the ICL setting with semantically unrelated labels, highlighting that sufficiently large models can perform linear classification tasks under this setting. In addition, they evaluate models fine-tuned with instructions and find that instruction tuning notably enhanced the utilization of semantic priors compared to learning input-label mappings from contextual demonstrations. Another influential work is the study by Wang et al. (2023) on the mechanism of ICL from the perspective of information flow, which find that in input-label pairs, label tokens act as anchors. Initially, the semantic information from the context aggregates into the token representation of the label tokens at the shallower layers of LLMs, and then the final predictions of LLMs reference the aggregated information in the label tokens. Building on this finding, anchor-based re-weighting methods, demonstration compression techniques, and diagnostic analysis frameworks for ICL errors are further proposed, yielding the expected results and validating the analysis. More recently, Cheng et al. (2025a) also argue from an empirical perspective and highlight that ICL fails to benefit reasoning models with a long chain-of-thought, reminding us that research on ICL needs to be considered from more multidimensional perspectives.
@@ 6.2.3 Inference-Time Scaling
Inference-time scaling represents a fundamental shift in the deployment of LLMs, where the reasoning capacity is no longer viewed as a static property of the model’s parameters but as a dynamic function of the computational resources allocated during interaction (Chen et al., 2025b; Snell et al., 2024a). This paradigm is primarily established through the Chain-of-Thought (CoT) mechanism and various external search-based algorithms that extend the model’s “thinking” process (Wei et al., 2022d; Yao et al., 2024a; Kang et al., 2024; Zhang et al., 2024a; Feng et al., 2023b). Based on current research, we categorize the theoretical investigations of this phenomenon into the following three core dimensions.
@@
Theoretical Expressivity and Boundaries of CoT. A foundational line of inquiry examines how the introduction of intermediate reasoning steps alters the inherent computational limits of the Transformer architecture. Theoretical analysis suggests that CoT serves as an effective depth-extender for auto-regressive models. Feng et al. (2023a) utilize circuit complexity theory to prove that finite-depth Transformers can perfectly execute these tasks by extending their effective depth linearly with the number of generated reasoning steps. This is further formalized by Li et al. (2024b), which demonstrates that while constant-depth Transformers without CoT are restricted to parallelizable complexity classes such as $AC^{0}$ or $TC^{0}$, the addition of reasoning steps enables the model to solve any problem within the $P/poly$ complexity class. To bridge these theoretical findings with practical performance, Chen et al. (2024c) introduce the Reasoning Boundary Framework (RBF) to define the quantitative limits of model performance across different task complexities. Furtherly, Sprague et al. (2025) reveal that CoT benefits are predominantly concentrated in mathematical and symbolic tasks, providing minimal gains in general knowledge retrieval or tasks lacking explicit logical operators.
@@
Mechanistic Origins and Internal Dynamics. Understanding how these reasoning capabilities emerge and are organized internally is critical for piercing the “black box” of LLM intelligence. Dutta et al. (2024) identify a functional bifurcation within the Transformer layers: lower layers primarily transform representations from pre-training priors to context-aware embeddings, while middle-to-higher layers act as answer writers that causally integrate information from previously generated CoT steps. Alternatively, Wang et al. (2024b) discover that reasoning circuits only form through “grokking”, when training significantly beyond the point of overfitting, allowing for robust out-of-distribution generalization that shallow statistical matching cannot achieve. Furthermore, Li et al. (2025a) provide a convergence analysis for how gradient descent optimization enables non-linear Transformers to learn CoT, quantifying the sample complexity required to remain robust against noisy context examples. Beyond explicit prompting, Wang and Zhou (2024) demonstrate that reasoning trajectories are an intrinsic capability of pre-trained models that can be elicited through specialized decoding strategies, such as exploring top-$k$ alternative tokens to find valid reasoning paths without human instructions.
@@
Reliability & Generalization Limits. Despite the empirical success of inference-time scaling, researchers have identified significant bottlenecks related to error accumulation and distributional sensitivity. Zhao et al. (2025a) argue that the efficacy of CoT is inherently fragile, relying heavily on the consistency between the training reasoning paths and the test-time queries, which suggests that models may be performing advanced pattern matching rather than deep logical deduction. This fragility is particularly evident in complex environments. Stechly et al. (2024) show that CoT performance degrades rapidly when task scale or complexity exceeds the scope of the provided examples. To mitigate the “snowball error” effect, where a single early mistake leads to catastrophic reasoning failure, Gan et al. (2025a) demonstrate that external scaling through search algorithms like Best-of-N and Monte Carlo Tree Search (MCTS) effectively expands the solution space and allows verifiers to select correct paths. To evaluate the quality of these dynamic steps, Ton et al. (2025) propose an “Information-Gain” metric, identifying “thinking tokens” that significantly reduce the predictive cross-entropy loss of the final answer, thereby providing a principled tool for diagnosing and optimizing the reasoning process.
@@ 6.3.1 The Overthinking Phenomenon
The empirical success of inference-time scaling, exemplified by leading reasoning models (OpenAI, 2024; Guo et al., 2025), has popularized the paradigm that more computation leads to better reasoning. However, recent research has identified a critical counter-phenomenon known as “overthinking”, where models generate excessive, redundant, or even erroneous reasoning steps for tasks that are either inherently simple or unsolvable.
Traditional intuition suggests that longer CoT sequences facilitate better task decomposition. However, Wu et al. (2025d) challenge this assumption by demonstrating an inverted U-shaped relationship between reasoning length and accuracy. This work posits that while length helps reduce sub-task difficulty, it simultaneously increases the risk of error accumulation. This balance is further discussed by Gan et al. (2025b), which treats CoT as an optimization process in continuous semantic space, identifying a fundamental trade-off between “under-reasoning” (underfitting) and “overthinking” (overfitting).
Overthinking is particularly prevalent when models mimic long-reasoning behaviors for trivial queries. Chen et al. (2025c) observe that models often generate verbose reasoning for extremely simple arithmetic, significantly increasing latency and cost without any performance gain. This dilemma is also explored in agentic contexts, Cuadron et al. (2025) highlight how excessive searching and value modeling can hinder the efficiency of logical agents. A more severe form of overthinking occurs when models encounter pathological queries. Fan et al. (2025) find that models optimized for reasoning tend to fall into redundant loops of self-doubt and hallucination when faced with unsolvable problems due to missing premises. This behavior is attributed to current RL mechanisms that over-reward detailed CoT. To address these inefficiencies, Li et al. (2025h) propose information-theoretic metrics to quantify the information contribution of each reasoning step. Their findings suggest that a significant portion of steps in modern reasoning models are low-entropy redundancies that can be compressed without compromising accuracy.
@@ 6.3.2 Latent Reasoning
Recent research explores bypassing these limitations by conducting reasoning directly within the model’s latent activation space. Hao et al. (2025) propose COCONUT, which allows models to reason in a continuous latent space by feeding hidden states back as subsequent inputs, enabling the encoding of multiple reasoning branches simultaneously. The architectural backbone of latent reasoning often involves weight-tied recurrence. Saunshi et al. (2025a) posit that reasoning performance is primarily driven by computational depth rather than total parameters. The study demonstrates that looped architectures can simulate CoT internally through “latent thoughts”. These models show strong inductive biases for reasoning tasks, suggesting that latent iterations can efficiently substitute for explicit token generation.
A pivotal development in this area is the study of the superposition mechanism within continuous latent spaces. Zhu et al. (2025b) demonstrate that the model can maintain multiple reasoning trajectories in a state of superposition within the continuous latent space, facilitating implicit parallel thinking that far exceeds the capabilities of traditional serial reasoning. The emergence of this mechanism is deeply tied to the training dynamics. Zhu et al. (2025a) further characterize it as a two-stage process and elucidate how the model can simultaneously maintain multiple inference traces in a continuous latent space, thereby achieving implicit parallel thinking. Despite its efficiency, latent reasoning introduces unique challenges. Xu and Sato (2025) highlight that while latent thoughts support more efficient parallel computation, discrete CoT remains superior for tasks requiring stochastic decoding to approximate complex solutions. More recently, Zou et al. (2026b) theoretically characterize latent reasoning and prove that high certainty enables precise execution but inhibits exploration. This work formalizes the capabilities and fundamental limits of latent CoT.
@@ 7.1 Fundamental Problems
The evaluation deeply intertwines with theoretical questions of metrology and security. Unlike in traditional machine learning, where concepts like “robustness”, “fairness”, and “privacy” were often well-defined and could be formalized using precise mathematical objectives and metrics, the current landscape of LLMs presents a new challenge (Chang et al., 2024; Anwar et al., 2024; Dominguez-Olmedo et al., 2025; Hardt and Mendler-Dünner, 2025). The core, fundamental problems in the Evaluation Stage are therefore:
@@
These two fundamental problems define the ultimate theoretical challenges in the Evaluation Stage. In fig.˜7 we provide a landscape of the corresponding topics. To begin answering these profound questions, the academic community has initiated several concrete lines of research, each tackling a specific evaluation tool or observable phenomenon. In what follows, we will review these research efforts, detailing how the study of specific strategies provides valuable insights.
@@ 7.2.1 Benchmark Theory
Benchmark Validity. A growing body of theoretical and empirical work suggests that high performance on static benchmarks may not correlate with true, generalized capabilities (Zhang et al., 2025a). Several key limitations in current benchmark-based evaluations are identified. One primary issue is “Shortcut Learning” (Du et al., 2023), where models are found to rely on spurious, non-robust features or biases within a dataset rather than mastering the high-level semantic or reasoning skills the benchmark purports to measure (Chandak et al., 2025). This leads to a fundamental lack of robustness and exposes that models may be overfitting to the specific artifacts of the test set rather than the underlying task (Lunardi et al., 2025). Furthermore, many traditional benchmarks are facing issues of saturation, where top-tier models approach perfect scores, limiting the benchmark’s ability to distinguish between SOTA models. To address this, Zhou et al. (2025a) have proposed applying frameworks from psychometrics, such as Item Response Theory, to analyze benchmark quality. This has revealed that many benchmarks suffer from a low “difficulty ceiling” and “item saturation”. Zhang and Hardt (2024) further derive analysis from the perspective of social choice theory, and demonstrate the trade-offs between benchmark diversity and stability. Finally, the single-scalar scores produced by most benchmarks obscure the complex combination of skills required for a task. Kim et al. (2025a) aim to mechanistically diagnose benchmark composition, decomposing performance into contributions from discrete cognitive abilities.
@@
LLM-as-a-Judge. To overcome the limitations of static benchmarks, especially for evaluating open-ended generation, the “LLM-as-a-Judge” (LLM-Judges) paradigm has become widespread (Gu et al., 2025). This approach leverages a powerful LLM to score or rank the outputs of other models. However, this has shifted the theoretical burden from evaluating task performance to evaluating the evaluator itself. This paradigm rests on several core assumptions that LLMs can serve as valid human proxies, are capable evaluators, are scalable, and are cost-effective, unfortunately, all of which are being theoretically challenged (Dorner et al., 2025). Researchers have argued that LLM-Judges may possess only “face validity” rather than true, robust evaluative capacity (Chehbouni et al., 2025). Two critical flaws being investigated are low reliability and poor psychometric validity. The common practice of using fixed randomness to ensure reproducibility does not guarantee internal consistency. By applying psychometric measures like McDonald’s omega and repeating evaluations with different random seeds, studies have found the internal consistency reliability of LLM-Judges to be questionable (Schroeder and Wood-Doughty, 2024). Also, LLM-Judges suffer from severe design flaws that can render their judgments noisy. Key issues include low “schematic adherence” and “factor collapse” causing the misalignment between the evaluation results and criteria (Feuer et al., 2025). Finally, LLM-Judges have been shown to exhibit numerous systematic biases, including position bias, verbosity bias, and authority bias (Ye et al., 2024b; Chen et al., 2024a). While some work suggests these biases can be partially mitigated through robust prompting with detailed scoring rubrics (Gao et al., 2025), the theoretical understanding of these biases remains a critical open area.